Translate the problem into a pair of linear equations in two variables. Solve the equations using either elimination or substitution. State your answer for the specified variable.
A child's bank contains $2.24 in pennies and nickels. If the number of pennies is 40 less than 3 times the number of nickels, then how many pennies are in the bank?

Answer by dabanfield(803)

Translate the problem into a pair of linear equations in two variables. Solve the equations using either elimination or substitution. State your answer for the specified variable.
A child's bank contains $2.24 in pennies and nickels. If the number of pennies is 40 less than 3 times the number of nickels, then how many pennies are in the bank?
Let x be the number of nickels and y the number of pennies. We have then:
1.) y = 3*x - 40
Since each nickle is worth .05 and each penny .01 we also have:
2.) .05*x + .01*y = 2.24
Multiplying both sides by 100 we have:
5x + y = 224
From the above we have y = -5x + 224
We can substitute -5x + 224 for y in equation 1.) and solve for x.
Once we have the value of x we can find y = -5x + 224.